# Maël Dumas LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Anthony Perez LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Mathis Rocton Algorithms and Complexity Group, TU Wien, Vienna, Austria Ioan Todinca LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France

### v3

@@ 4 Kernelization algorithm for Strictly Chordal Editing
Let us consider a positive instance $(G=(V,E),k)$ of Strictly Chordal Editing, $F$ a suitable solution and $H=G\triangle F$. Denote by $\mathcal{C}(H)$ the critical clique graph of $H$ as described in Definition 2. Since $|F|\leqslant k$, we know that at most $2k$ critical cliques of $\mathcal{C}(H)$ may contain affected vertices. Let $A$ be the set of such critical cliques, $T$ the minimum subgraph of $\mathcal{C}(H)$ that spans all critical cliques of $A$ and $A^{\prime}$ the set of critical cliques of degree at least $3$ in $T$. Recall that $|A^{\prime}\setminus A|\leqslant 3\cdot|A|$ (Lemma 5). We will define the notion of SC-branch, corresponding to well-structured strictly chordal subgraphs of $G$. We will focus our interest on two types of SC-branches: the ones that are connected to the rest of the graph by only one critical clique, called $1$-SC-branches, and the ones that are connected to the rest of the graph by exactly two critical cliques, called $2$-SC-branches. This time the $1$-SC-branches will only be reduced to two critical cliques. The connected components of the graph $T\setminus(A\cup A^{\prime})$ are paths, which correspond to subgraphs of $2$-SC-branches, and we will bound their length by $k+3$. We know that there are at most $4k$ such connected components, thus there is $O(k^{2})$ critical cliques in $T$. Finally, the connected components of the graph $\mathcal{C}(H)\setminus V(T)$ correspond to $1$-SC-branches or sets of connected $1$-SC-branches. The latter case will require a specific reduction rule to be dealt with. This implies that every critical clique or maximal clique of $T$ can have a linear number of $1$-SC-branches of $\mathcal{C}(H)\setminus V(T)$ adjacent to it. Altogether, the graph $\mathcal{C}(H)$ contains $O(k^{3})$ critical cliques, and each critical clique is of size at most $k+1$, hence the graph $G$ contains $O(k^{4})$ vertices.
@@ 4.2 Reducing Strictly Chordal branches
Now suppose that $Q_{i}$, $1\leqslant i\leqslant r$ is neither a critical clique nor a maximal clique of $G_{i}$ and intersects at least two maximal cliques of $G_{i}$. Suppose first that $Q_{i}$ is not a module, then there exist $u,v\in Q_{i}$ and $x\in V(G_{i})\setminus Q_{i}$ adjacent to only $u$ or $v$. Since $Q_{i}$ is not a maximal clique of $G_{i}$, there exists $y\in V(G_{i})\setminus Q_{i}$ adjacent to $u$ and $v$. Together with $w\in P$, the vertex set $\{u,v,w,x,y\}$ induces in $B$ a gem if $x$ and $y$ are adjacent and a dart otherwise, leading to a contradiction. If $Q_{i}$ is a module, since it is not a critical clique (i.e. a maximal clique module) there exist $u\in Q_{i}$ and $v\in V(G_{i})\setminus Q_{i}$ that lie in the same critical clique of $G_{i}$. Since $Q_{i}$ intersects two maximal cliques of $G_{i}$, there exist two non-adjacent vertices $x,y\in V(G_{i})\setminus Q_{i}$ that are adjacent to both $u$ and $v$. Together with $w\in P$, the vertex set $\{u,v,w,x,y\}$ induces a dart in $B$, leading to a contradiction. It remains that $Q_{i}$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $G_{i}$. ∎
Lemma 16.
@@
Proof.
Suppose that $C^{\prime}$ is not a critical clique of $H^{\prime}$ and intersects at least two maximal cliques of $H^{\prime}$. Since $C^{\prime}\subseteq C$ is a module but not a critical clique, it is not a maximal clique module and thus there exist $u\in C^{\prime}$ and $v\notin C^{\prime}$ that lie in the same critical clique in $H^{\prime}$. Notice that there must exists a vertex $w\in N_{H}(C^{\prime})\cap V(B^{R})$ adjacent to exactly one of $u$ and $v$ in $H$, since otherwise we would have $N_{H}(u)\cap V(B)=N_{H}(v)\cap V(B)$ and $N_{H^{\prime}}(u)=N_{H^{\prime}}(v)$, meaning that $C$ is not a critical clique of $H$, a contradiction. Since $C^{\prime}$ intersects two maximal cliques of $H^{\prime}$ there exist two non-adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertex set $\{u,v,w,x,y\}$ induces a dart in $H$ if $w$ is not adjacent to neither $x$ nor $y$, a gem if it is adjacent to one of them, and if $w$ is adjacent to both $x$ and $y$, $\{w,x,v,y\}$ induces a $C_{4}$ in $H$. It remains that $C^{\prime}$ is a critical clique of $H^{\prime}$ or $C^{\prime}$ is included in exactly one maximal clique of $H^{\prime}$ if $N_{H}(C^{\prime})\cap V(B^{R})\neq\emptyset$. ∎
@@
If $P\cap N=\emptyset$, let $H^{*}$ be the strictly chordal graph corresponding to the disjoint union of $H^{\prime}$ and the graphs $G_{1},\dots,G_{r}$. Let $F^{*}$ be such that $H^{*}=G\triangle F^{*}$, in both cases by construction $|F^{*}|\leqslant|F|$ and the desired properties are verified. ∎
Figure 7: Illustration of the proof of Lemma 17 for $r=3$. The graph $H^{*}=(G_{1}\cup G_{2}\cup G_{3},Q_{1}\cup Q_{2}\cup Q_{3})\otimes(H^{\prime},C^{\prime})$ is strictly chordal by Lemma 4.
Rule 4.3.
@@
Proof.
Let $G^{\prime}$ be the graph obtained from an application of Rule 4.3 on $1$-SC-branch $B$ of graph $G$ with attachment point $P$. We denote by $K$ be the clique that replaced $B$. Let $F$ be a $k$-edition of $G$ and $H=G\triangle F$. By Lemma 17 we can assume that $F$ only affects vertices of $P\cup N_{B}(P)$ and all vertices of $N_{B}(P)$ are adjacent to the same vertices of $V(G)\setminus V(B^{R})$. This implies that if $|N_{B}(P)|>k$ no $k$-edition can affect these vertices. Let $C$ be the critical clique of $H$ which contains $P$ and $C^{\prime}=C\setminus V(B^{R})$. By Lemma 16, $C^{\prime}$ is a critical clique or intersects exactly one maximal clique of $H_{a}=H\setminus V(B^{R})$. If $N_{B}(P)$ is adjacent to $C^{\prime}$ in $H$, consider the graph $H^{\prime}=(H_{a},C^{\prime})\otimes(K,K)$. Since $K$ is clique, by Lemma 4, $H^{\prime}$ is strictly chordal. Else, consider the strictly chordal graph $H^{\prime}$ corresponding to the disjoint union of $H_{a}$ and $K$. Let $F^{\prime}$ be the edition such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$. Since $|K|=min\{|N_{B}(P)|,k+1\}$, by construction we have $|F|=|F^{\prime}|$, hence $F^{\prime}$ is a $k$-edition of $G^{\prime}$.
@@
If the clique $K$ is adjacent to $C^{\prime}$ in $H^{\prime}$, consider the graph $H=(H_{a}^{\prime},C^{\prime})\otimes(\bigcup_{1\leqslant i\leqslant r}G_{i},\bigcup_{1\leqslant i\leqslant r}Q_{i})$. By Lemma 4 and Observation 1, $H^{\prime}$ is strictly chordal. Else, consider the strictly chordal graph $H$ corresponding to the disjoint union of $H_{a}^{\prime}$ and $G_{i}$ for $1\leqslant i\leqslant r$. Let $F$ be the edition such that $H=G\triangle F$, since $|K|=min\{|N_{B}(P)|,k+1\}$, by construction we have $|F|=|F^{\prime}|$, hence $F$ is a $k$-edition of $G$. ∎
@@
Finally, since $N$ is a module in $H^{\prime}$, if it is not a critical clique, there exist $u\in N$, a vertex $v$ in the same critical clique as $u$ in $H^{\prime}$ and a vertex $w\in\bigcup_{i=1}^{l}P_{i}$ not affected by $F$ such that $vw\notin E(H)$. If $N$ intersects at least two maximal cliques of $H^{\prime}$ there exist two non-adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertex set $\{u,v,w,x,y\}$ induces a dart in $H$ since $w$ is only adjacent to $u$. It remains that $N$ is a critical clique or intersects exactly one maximal clique of $H^{\prime}$. ∎
Rule 4.4.
@@
Proof.
Let $G^{\prime}$ be the graph obtained after the application of Rule 4.4 on $B_{1},\dots,B_{l}$, and $S$ the new vertices introduced by this rule. Notice that the subgraph $G_{S}^{\prime}=G^{\prime}[S]$ is strictly chordal and form two $1$-SC-branches that satisfy the conditions of Lemma 19. First, let $F$ be a $k$-edition of $G$, $H=G\triangle F$ and $H_{a}=H\setminus\bigcup_{i=1}^{l}V(B_{i})$. By Lemma 19 $N$ is a critical clique, a maximal clique (unless $r>1$) or intersects exactly one maximal clique of $H_{a}$. Therefore, by Lemma 4 and Observation 1 the graph $H^{\prime}=(H_{a},N)\otimes(G_{S}^{\prime},K\cup K^{\prime})$ is strictly chordal and the edition $F^{\prime}$ such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$ is of size at most $|F|\leqslant k$.
Conversely, let $F^{\prime}$ be a $k$-edition of $G^{\prime}$, $H^{\prime}=G^{\prime}\triangle F^{\prime}$ and $H_{a}^{\prime}=H^{\prime}\setminus S$. Notice that the subgraph $G_{b}=G[\bigcup_{i=1}^{l}V(B_{i})]$ is strictly chordal. By Lemma 19 $N$ is a critical clique, a maximal clique (unless $r>1$) or intersects exactly one maximal clique of $H_{a}^{\prime}$. Let $Q_{1},\dots,Q_{r}$ the cliques of $G_{b}$ adjacent to $N$, using arguments similar to the ones of Lemma 15, it is easy to see that each $Q_{i}$ is critical clique, a maximal clique or intersects exactly one maximal clique of $G_{b}$. Therefore, by Lemma 4 and Observation 1 the graph $H=(H_{a}^{\prime},N)\otimes(G_{b},Q_{1}\cup\dots\cup Q_{r})$ is strictly chordal and the edition $F$ such that $H=G\triangle F$ is of size at most $|F^{\prime}|\leqslant k$. ∎
@@ 4.3 Reducing the 2-SC-branches
We first consider the case where $P_{1}$ and $P_{2}$ are in different connected component of $H^{\prime}$. Let $H^{\prime}_{1}$ and $H^{\prime}_{2}$ be the two connected components of $H^{\prime}$ containing $P_{1}$ and $P_{2}$, respectively, and $H^{\prime}_{3}$ the remaining of $H^{\prime}$. For $i\in\{1,2\}$, Lemma 15 implies that $N_{B}(P_{i})$ is a clique in $G$ since $B^{R}$ is connected, and $N_{B}(P_{i})$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $B^{R}$. By Lemma 16, $C_{i}^{\prime}$ is a critical clique or intersects exactly one maximal clique of the connected component $H^{\prime}_{i}$.
By Lemma 4, the graph $H^{*}$ corresponding to the disjoint union of
@@
Figure 8: Illustration of the case where $B\triangle F$ is connected in Lemma 21.
@@
Hence we can assume that $F$ contains an edge-cut of $B$. First, if $F_{1}=(P_{1}\times N_{B}(P_{1}))\subseteq F$, consider the graph $G_{1}=G\triangle F_{1}$. Observe that $B_{1}=B\setminus P_{1}$ is a $1$-SC-branch of $G_{1}$ with attachment point $P_{2}$. By Lemma 17 there exists an optimal edition $F_{2}$ of $G_{1}$ where the vertices of $B_{1}$ affected by $F_{2}$ are in $P_{2}\cup N_{B_{1}}(P_{2})$. By Observation 2, $F_{1}\cup F_{2}$ is an optimal edition and it verifies the desired properties. The same arguments can be used if $(P_{2}\times N_{B}(P_{2}))\subseteq F$.
Now, if $F$ contains neither $(P_{1}\times N_{B}(P_{1}))$ nor $(P_{2}\times N_{B}(P_{2}))$, let $F_{1}=F\cap(V(B^{R})\times V(B^{R}))$. There are two components in $B\triangle F_{1}$, say $B_{1}$ the one containing $P_{1}$ and $B_{2}$ one containing $P_{2}$. These connected components are $1$-SC-branches of $G_{1}=G\triangle F_{1}$ with attachment points $P_{1}$ and $P_{2}$. As before Lemma 17 applies on the $1$-SC-branches $B_{1}$ and $B_{2}$, thus there is an optimal edition $F_{2}$ of $G_{1}$ where the only vertices of $B_{1}$ and $B_{2}$ affected are included in $P_{1}\cup N_{B}(P_{1})$ and $P_{2}\cup N_{B}(P_{2})$. If $F_{1}$ is a min-cut of $B$, the edition $F_{1}\cup F_{2}$ is an optimal edition of $G$ that verifies the desired properties, else, replacing $F_{1}$ by a min-cut of $B$ results in a smaller edition that verifies the properties. ∎
Rule 4.5.
@@
Let $F$ be a $k$-edition of $G$ that satisfies Lemma 21 and $H=G\triangle F$. Observe that if $mc(B)\geqslant k+1$, then $F$ does not affect vertices of $V(B)$ other than $P_{1}\cup N_{B}(P_{1})\cup P_{2}\cup N_{B}(P_{2})$ and if $|N_{B}(P_{i})|>k,i\in\{1,2\}$ $F$ does not affect the vertices of $N_{B}(P_{i})$. Let $C_{i}$ be the critical clique of $H$ which contains $P_{i}$ and $C_{i}^{\prime}=C_{i}\setminus B^{R}$. By Lemma 16, $C^{\prime}_{i}$ is a critical clique or intersects exactly one maximal clique of $H_{a}=H\setminus V(B^{R})$. Observe moreover that $N_{B^{\prime}}(P_{i})$ is critical clique of $B^{\prime R}$.
We now construct a $k$-edition $F^{\prime}$ of $G^{\prime}$. We consider first the tree cases where $F$ contains an edge-cut of $B$. If w.l.o.g. $(P_{1}\times N_{B}(P_{1}))\subseteq F$ and $(P_{2}\times N_{B}(P_{2}))\cap F=\emptyset$. Observe that $(C_{1}^{\prime}\times N_{B^{\prime}}(P_{1})\subseteq F$ since $C_{1}$ is a critical clique of $H$. By Lemma 4 the graph $H^{\prime}=((H_{a},C_{1}^{\prime})\otimes(B^{\prime R},N_{B^{\prime}}(P_{1}))$ is strictly chordal. If $(P_{1}\times N_{B}(P_{1}))\subseteq F$ and $(P_{2}\times N_{B}(P_{2}))\subseteq F$ then $H^{\prime}$ the graph resulting of the disjoint union of $H_{a}$ and $B^{\prime R}$ is strictly chordal. If $F$ contains a min-cut of $B$, then consider $B_{i}^{\prime},i\in\{1,2\}$ the connected component of $B^{\prime}\triangle F_{1}^{\prime}$ containing $P_{i}$, where $F_{1}^{\prime}$ is a min-cut of $B^{\prime}$. By Lemma 4 the graph $H^{\prime}=((H_{a},C_{1}^{\prime})\otimes(B_{1}^{\prime},N_{B^{\prime}}(P_{1})),C_{2}^{\prime})\otimes(B_{2}^{\prime},N_{B^{\prime}}(P_{2}))$ is strictly chordal.
Otherwise, $P_{1}$ and $P_{2}$ are not in the same connected component of $H_{a}$ by Lemma 21. Let $H_{1}$ and $H_{2}$ be these components and $H_{3}$ be the remaining of $H_{a}$. By Lemma 4, the graph $H^{\prime}$ corresponding to the disjoint union of $((B^{\prime R},N_{B^{\prime}}(P_{1}))\otimes(H_{1},C_{1}^{\prime}),N_{B^{\prime}}(P_{2}))\otimes(H_{2},C_{2}^{\prime})$ and $H_{3}$ is strictly chordal.

### v4

@@ 4 Kernelization algorithm for Strictly Chordal Editing
Let us consider a positive instance $(G=(V,E),k)$ of Strictly Chordal Editing, $F$ a suitable solution and $H=G\triangle F$. Denote by $\mathcal{C}(H)$ the critical clique graph of $H$ as described in Definition 2. Since $|F|\leqslant k$, we know that at most $2k$ critical cliques of $\mathcal{C}(H)$ may contain affected vertices. Let $A$ be the set of such critical cliques, $T$ the minimum subgraph of $\mathcal{C}(H)$ that spans all critical cliques of $A$ and $A^{\prime}$ the set of critical cliques of degree at least $3$ in $T$. Recall that $|A^{\prime}\setminus A|\leqslant 3\cdot|A|$ (Lemma 2.10). We will define the notion of SC-branch, corresponding to well-structured strictly chordal subgraphs of $G$. We will focus our interest on two types of SC-branches: the ones that are connected to the rest of the graph by only one critical clique, called $1$-SC-branches, and the ones that are connected to the rest of the graph by exactly two critical cliques, called $2$-SC-branches. This time the $1$-SC-branches will only be reduced to two critical cliques. The connected components of the graph $T\setminus(A\cup A^{\prime})$ are paths, which correspond to subgraphs of $2$-SC-branches, and we will bound their length by $k+3$. We know that there are at most $4k$ such connected components, thus there is $O(k^{2})$ critical cliques in $T$. Finally, the connected components of the graph $\mathcal{C}(H)\setminus V(T)$ correspond to $1$-SC-branches or sets of connected $1$-SC-branches. The latter case will require a specific reduction rule to be dealt with. This implies that every critical clique or maximal clique of $T$ can have a linear number of $1$-SC-branches of $\mathcal{C}(H)\setminus V(T)$ adjacent to it. Altogether, the graph $\mathcal{C}(H)$ contains $O(k^{3})$ critical cliques, and each critical clique is of size at most $k+1$, hence the graph $G$ contains $O(k^{4})$ vertices.
@@ 4.2 Reducing Strictly Chordal branches
Now suppose that $Q_{i}$, $1\leqslant i\leqslant r$ is neither a critical clique nor a maximal clique of $G_{i}$ and intersects at least two maximal cliques of $G_{i}$. Suppose first that $Q_{i}$ is not a module, then there exist $u,v\in Q_{i}$ and $x\in V(G_{i})\setminus Q_{i}$ adjacent to only $u$ or $v$. Since $Q_{i}$ is not a maximal clique of $G_{i}$, there exists $y\in V(G_{i})\setminus Q_{i}$ adjacent to $u$ and $v$. Together with $w\in P$, the vertex set $\{u,v,w,x,y\}$ induces in $B$ a gem if $x$ and $y$ are adjacent and a dart otherwise, leading to a contradiction. If $Q_{i}$ is a module, since it is not a critical clique (i.e. a maximal clique module) there exist $u\in Q_{i}$ and $v\in V(G_{i})\setminus Q_{i}$ that lie in the same critical clique of $G_{i}$. Since $Q_{i}$ intersects two maximal cliques of $G_{i}$, there exist two non-adjacent vertices $x,y\in V(G_{i})\setminus Q_{i}$ that are adjacent to both $u$ and $v$. Together with $w\in P$, the vertex set $\{u,v,w,x,y\}$ induces a dart in $B$, leading to a contradiction. It remains that $Q_{i}$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $G_{i}$.
Lemma 4.32.
@@
Proof 4.33.
Suppose that $C^{\prime}$ is not a critical clique of $H^{\prime}$ and intersects at least two maximal cliques of $H^{\prime}$. Since $C^{\prime}\subseteq C$ is a module but not a critical clique, it is not a maximal clique module and thus there exist $u\in C^{\prime}$ and $v\notin C^{\prime}$ that lie in the same critical clique in $H^{\prime}$. Notice that there must exists a vertex $w\in N_{H}(C^{\prime})\cap V(B^{R})$ adjacent to exactly one of $u$ and $v$ in $H$, since otherwise we would have $N_{H}(u)\cap V(B)=N_{H}(v)\cap V(B)$ and $N_{H^{\prime}}(u)=N_{H^{\prime}}(v)$, meaning that $C$ is not a critical clique of $H$, a contradiction. Since $C^{\prime}$ intersects two maximal cliques of $H^{\prime}$ there exist two non-adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertex set $\{u,v,w,x,y\}$ induces a dart in $H$ if $w$ is not adjacent to neither $x$ nor $y$, a gem if it is adjacent to one of them, and if $w$ is adjacent to both $x$ and $y$, $\{w,x,v,y\}$ induces a $C_{4}$ in $H$. It remains that $C^{\prime}$ is a critical clique of $H^{\prime}$ or $C^{\prime}$ is included in exactly one maximal clique of $H^{\prime}$ if $N_{H}(C^{\prime})\cap V(B^{R})\neq\emptyset$.
@@
If $P\cap N=\emptyset$, let $H^{*}$ be the strictly chordal graph corresponding to the disjoint union of $H^{\prime}$ and the graphs $G_{1},\dots,G_{r}$. Let $F^{*}$ be such that $H^{*}=G\triangle F^{*}$, in both cases by construction $|F^{*}|\leqslant|F|$ and the desired properties are verified.
Figure 7: Illustration of the proof of Lemma 4.34 for $r=3$. The graph $H^{*}=(G_{1}\cup G_{2}\cup G_{3},Q_{1}\cup Q_{2}\cup Q_{3})\otimes(H^{\prime},C^{\prime})$ is strictly chordal by Lemma 2.8.
Rule 4.3
@@
Proof 4.37.
Let $G^{\prime}$ be the graph obtained from an application of Rule 4.3 on $1$-SC-branch $B$ of graph $G$ with attachment point $P$. We denote by $K$ be the clique that replaced $B$. Let $F$ be a $k$-edition of $G$ and $H=G\triangle F$. By Lemma 4.34 we can assume that $F$ only affects vertices of $P\cup N_{B}(P)$ and all vertices of $N_{B}(P)$ are adjacent to the same vertices of $V(G)\setminus V(B^{R})$. This implies that if $|N_{B}(P)|>k$ no $k$-edition can affect these vertices. Let $C$ be the critical clique of $H$ which contains $P$ and $C^{\prime}=C\setminus V(B^{R})$. By Lemma 4.32, $C^{\prime}$ is a critical clique or intersects exactly one maximal clique of $H_{a}=H\setminus V(B^{R})$. If $N_{B}(P)$ is adjacent to $C^{\prime}$ in $H$, consider the graph $H^{\prime}=(H_{a},C^{\prime})\otimes(K,K)$. Since $K$ is clique, by Lemma 2.8, $H^{\prime}$ is strictly chordal. Else, consider the strictly chordal graph $H^{\prime}$ corresponding to the disjoint union of $H_{a}$ and $K$. Let $F^{\prime}$ be the edition such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$. Since $|K|=min\{|N_{B}(P)|,k+1\}$, by construction we have $|F|=|F^{\prime}|$, hence $F^{\prime}$ is a $k$-edition of $G^{\prime}$.
@@
If the clique $K$ is adjacent to $C^{\prime}$ in $H^{\prime}$, consider the graph $H=(H_{a}^{\prime},C^{\prime})\otimes(\bigcup_{1\leqslant i\leqslant r}G_{i},\bigcup_{1\leqslant i\leqslant r}Q_{i})$. By Lemma 2.8 and Observation 1, $H^{\prime}$ is strictly chordal. Else, consider the strictly chordal graph $H$ corresponding to the disjoint union of $H_{a}^{\prime}$ and $G_{i}$ for $1\leqslant i\leqslant r$. Let $F$ be the edition such that $H=G\triangle F$, since $|K|=min\{|N_{B}(P)|,k+1\}$, by construction we have $|F|=|F^{\prime}|$, hence $F$ is a $k$-edition of $G$.
@@
Finally, since $N$ is a module in $H^{\prime}$, if it is not a critical clique, there exist $u\in N$, a vertex $v$ in the same critical clique as $u$ in $H^{\prime}$ and a vertex $w\in\bigcup_{i=1}^{l}P_{i}$ not affected by $F$ such that $vw\notin E(H)$. If $N$ intersects at least two maximal cliques of $H^{\prime}$ there exist two non-adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertex set $\{u,v,w,x,y\}$ induces a dart in $H$ since $w$ is only adjacent to $u$. It remains that $N$ is a critical clique or intersects exactly one maximal clique of $H^{\prime}$.
Rule 4.4
@@
Proof 4.41.
Let $G^{\prime}$ be the graph obtained after the application of Rule 4.4 on $B_{1},\dots,B_{l}$, and $S$ the new vertices introduced by this rule. Notice that the subgraph $G_{S}^{\prime}=G^{\prime}[S]$ is strictly chordal and form two $1$-SC-branches that satisfy the conditions of Lemma 4.38. First, let $F$ be a $k$-edition of $G$, $H=G\triangle F$ and $H_{a}=H\setminus\bigcup_{i=1}^{l}V(B_{i})$. By Lemma 4.38 $N$ is a critical clique, a maximal clique (unless $r>1$) or intersects exactly one maximal clique of $H_{a}$. Therefore, by Lemma 2.8 and Observation 1 the graph $H^{\prime}=(H_{a},N)\otimes(G_{S}^{\prime},K\cup K^{\prime})$ is strictly chordal and the edition $F^{\prime}$ such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$ is of size at most $|F|\leqslant k$.
Conversely, let $F^{\prime}$ be a $k$-edition of $G^{\prime}$, $H^{\prime}=G^{\prime}\triangle F^{\prime}$ and $H_{a}^{\prime}=H^{\prime}\setminus S$. Notice that the subgraph $G_{b}=G[\bigcup_{i=1}^{l}V(B_{i})]$ is strictly chordal. By Lemma 4.38 $N$ is a critical clique, a maximal clique (unless $r>1$) or intersects exactly one maximal clique of $H_{a}^{\prime}$. Let $Q_{1},\dots,Q_{r}$ the cliques of $G_{b}$ adjacent to $N$, using arguments similar to the ones of Lemma 4.30, it is easy to see that each $Q_{i}$ is critical clique, a maximal clique or intersects exactly one maximal clique of $G_{b}$. Therefore, by Lemma 2.8 and Observation 1 the graph $H=(H_{a}^{\prime},N)\otimes(G_{b},Q_{1}\cup\dots\cup Q_{r})$ is strictly chordal and the edition $F$ such that $H=G\triangle F$ is of size at most $|F^{\prime}|\leqslant k$.
@@ 4.3 Reducing the 2-SC-branches
We first consider the case where $P_{1}$ and $P_{2}$ are in different connected component of $H^{\prime}$. Let $H^{\prime}_{1}$ and $H^{\prime}_{2}$ be the two connected components of $H^{\prime}$ containing $P_{1}$ and $P_{2}$, respectively, and $H^{\prime}_{3}$ the remaining of $H^{\prime}$. For $i\in\{1,2\}$, Lemma 4.30 implies that $N_{B}(P_{i})$ is a clique in $G$ since $B^{R}$ is connected, and $N_{B}(P_{i})$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $B^{R}$. By Lemma 4.32, $C_{i}^{\prime}$ is a critical clique or intersects exactly one maximal clique of the connected component $H^{\prime}_{i}$.
By Lemma 2.8, the graph $H^{*}$ corresponding to the disjoint union of
@@
Figure 8: Illustration of the case where $B\triangle F$ is connected in Lemma 4.42.
@@
Hence we can assume that $F$ contains an edge-cut of $B$. First, if $F_{1}=(P_{1}\times N_{B}(P_{1}))\subseteq F$, consider the graph $G_{1}=G\triangle F_{1}$. Observe that $B_{1}=B\setminus P_{1}$ is a $1$-SC-branch of $G_{1}$ with attachment point $P_{2}$. By Lemma 4.34 there exists an optimal edition $F_{2}$ of $G_{1}$ where the vertices of $B_{1}$ affected by $F_{2}$ are in $P_{2}\cup N_{B_{1}}(P_{2})$. By Observation 2, $F_{1}\cup F_{2}$ is an optimal edition and it verifies the desired properties. The same arguments can be used if $(P_{2}\times N_{B}(P_{2}))\subseteq F$.
Now, if $F$ contains neither $(P_{1}\times N_{B}(P_{1}))$ nor $(P_{2}\times N_{B}(P_{2}))$, let $F_{1}=F\cap(V(B^{R})\times V(B^{R}))$. There are two components in $B\triangle F_{1}$, say $B_{1}$ the one containing $P_{1}$ and $B_{2}$ one containing $P_{2}$. These connected components are $1$-SC-branches of $G_{1}=G\triangle F_{1}$ with attachment points $P_{1}$ and $P_{2}$. As before Lemma 4.34 applies on the $1$-SC-branches $B_{1}$ and $B_{2}$, thus there is an optimal edition $F_{2}$ of $G_{1}$ where the only vertices of $B_{1}$ and $B_{2}$ affected are included in $P_{1}\cup N_{B}(P_{1})$ and $P_{2}\cup N_{B}(P_{2})$. If $F_{1}$ is a min-cut of $B$, the edition $F_{1}\cup F_{2}$ is an optimal edition of $G$ that verifies the desired properties, else, replacing $F_{1}$ by a min-cut of $B$ results in a smaller edition that verifies the properties.
Rule 4.5
@@
Let $F$ be a $k$-edition of $G$ that satisfies Lemma 4.42 and $H=G\triangle F$. Observe that if $mc(B)\geqslant k+1$, then $F$ does not affect vertices of $V(B)$ other than $P_{1}\cup N_{B}(P_{1})\cup P_{2}\cup N_{B}(P_{2})$ and if $|N_{B}(P_{i})|>k,i\in\{1,2\}$ $F$ does not affect the vertices of $N_{B}(P_{i})$. Let $C_{i}$ be the critical clique of $H$ which contains $P_{i}$ and $C_{i}^{\prime}=C_{i}\setminus B^{R}$. By Lemma 4.32, $C^{\prime}_{i}$ is a critical clique or intersects exactly one maximal clique of $H_{a}=H\setminus V(B^{R})$. Observe moreover that $N_{B^{\prime}}(P_{i})$ is critical clique of $B^{\prime R}$.
We now construct a $k$-edition $F^{\prime}$ of $G^{\prime}$. We consider first the tree cases where $F$ contains an edge-cut of $B$. If w.l.o.g. $(P_{1}\times N_{B}(P_{1}))\subseteq F$ and $(P_{2}\times N_{B}(P_{2}))\cap F=\emptyset$. Observe that $(C_{1}^{\prime}\times N_{B^{\prime}}(P_{1})\subseteq F$ since $C_{1}$ is a critical clique of $H$. By Lemma 2.8 the graph $H^{\prime}=((H_{a},C_{1}^{\prime})\otimes(B^{\prime R},N_{B^{\prime}}(P_{1}))$ is strictly chordal. If $(P_{1}\times N_{B}(P_{1}))\subseteq F$ and $(P_{2}\times N_{B}(P_{2}))\subseteq F$ then $H^{\prime}$ the graph resulting of the disjoint union of $H_{a}$ and $B^{\prime R}$ is strictly chordal. If $F$ contains a min-cut of $B$, then consider $B_{i}^{\prime},i\in\{1,2\}$ the connected component of $B^{\prime}\triangle F_{1}^{\prime}$ containing $P_{i}$, where $F_{1}^{\prime}$ is a min-cut of $B^{\prime}$. By Lemma 2.8 the graph $H^{\prime}=((H_{a},C_{1}^{\prime})\otimes(B_{1}^{\prime},N_{B^{\prime}}(P_{1})),C_{2}^{\prime})\otimes(B_{2}^{\prime},N_{B^{\prime}}(P_{2}))$ is strictly chordal.
Otherwise, $P_{1}$ and $P_{2}$ are not in the same connected component of $H_{a}$ by Lemma 4.42. Let $H_{1}$ and $H_{2}$ be these components and $H_{3}$ be the remaining of $H_{a}$. By Lemma 2.8, the graph $H^{\prime}$ corresponding to the disjoint union of $((B^{\prime R},N_{B^{\prime}}(P_{1}))\otimes(H_{1},C_{1}^{\prime}),N_{B^{\prime}}(P_{2}))\otimes(H_{2},C_{2}^{\prime})$ and $H_{3}$ is strictly chordal.